# Deep learning-driven likelihood-free parameter inference for 21-cm forest observations

### v1

@@ Methods
Expected 1D power spectrum of the 21-cm forest. A total of 100 measurements of 10 comoving Mpc length segments along the line-of-sight neutral patch against 10 background light sources with $S_{150}$ = 10 mJy yielded the expected value for the 1D power spectrum of 21-cm forest at $z$ = 9. The left panels present the 1D power spectra for different $m_{\rm WDM}$ values in the WDM models. Blue, orange, and green correspond to $m_{\rm WDM}$ = 3 keV, $m_{\rm WDM}$ = 6 keV and $m_{\rm WDM}$ = 9 keV, respectively. $T_{\rm K}$ in the upper left panel are both $T_{\rm K}$ = 60 K, and $T_{\rm K}$ in the lower left panel are all $T_{\rm K}$ = 600 K. The right panel presents the 1D power spectra corresponding to the different $T_{\rm K}$. In all cases, $m_{\rm WDM}$ is set to 6 keV. The blue, orange, and green in the upper right panel correspond to $T_{\rm K}$ = 40 K, $T_{\rm K}$ = 60 K and $T_{\rm K}$ = 80 K. The blue, orange, and green in the lower right panel correspond to $T_{\rm K}$ = 400 K, $T_{\rm K}$ = 600 K and $T_{\rm K}$ = 800 K, respectively. The black solid and dashed lines are the thermal noises $P^{\mathrm{N}}$ expected for an integration time of 100 h using SKA1-LOW and an integration time of 200 h using SKA2-LOW, respectively. The error bars show sample variances of the 1D power spectra of the 21-cm forest.
@@
Constraints on $m_{\rm WDM}$ and $T_{\rm K}$ using different scales of 1D power spectra with the INF without considering thermal noise. The blue contour is based on the INF for $k$ in [$0.8~{}\rm Mpc^{-1}$, $3.5~{}\rm Mpc^{-1}$], the green contour is based on the INF for $k$ in [$5.5~{}\rm Mpc^{-1}$, $22.0~{}\rm Mpc^{-1}$], the orange contour is based on the INF for $k$ in [$35.2~{}\rm Mpc^{-1}$, $144.5~{}\rm Mpc^{-1}$] and the black contour is based on the INF for total scale. The panel presents the results when $m_{\rm WDM}$ = 6 keV and $T_{\rm K}$ = 60 K with an integration time of 100 h using SKA1-LOW.
@@
P-P plot of the INF parameter inference. The left panel presents the P-P plot of the INF inference with an integration time of 100 h using SKA1-LOW and for a total of 2700 power spectra with $m_{\rm WDM}$ ranging from 5 keV to 7 keV and $T_{\rm K}$ from 50 K to 70 K. The right panel presents the P-P plot of the INF inference with an integration time of 200 h using SKA2-LOW and for a total of 2700 power spectra with $m_{\rm WDM}$ ranging from 5 keV to 7 keV and $T_{\rm K}$ from 400 K to 800 K. The gray area is the range of the $3\sigma$ confidence interval. KS test $p$-values are denoted in the legend.
@@
Expected brightness temperature of the 21-cm forest. A total of 10 measurements of 10 comoving megaparsec length segments along a line-of-sight neutral patch against 1 background light source with $S_{150}$ = 10 mJy yielded the expected value for the brightness temperature of 21-cm forest at $z$ = 9. The left panels present the brightness temperature for different $m_{\rm WDM}$ values in the WDM models. Blue, orange, and green correspond to $m_{\rm WDM}$ = 3 keV, $m_{\rm WDM}$ = 6 keV and $m_{\rm WDM}$ = 9 keV, respectively. $T_{\rm K}$ in the upper left panel are both $T_{\rm K}$ = 60 K, and $T_{\rm K}$ in the lower left panel are all $T_{\rm K}$ = 600 K. The right panels present the brightness temperature corresponding to the different $T_{\rm K}$. In all cases, $m_{\rm WDM}$ is set to 6 keV. The blue, orange, and green in the upper right panel correspond to $T_{\rm K}$ = 40 K, $T_{\rm K}$ = 60 K and $T_{\rm K}$ = 80 K. The blue, orange, and green in the lower right panel correspond to $T_{\rm K}$ = 400 K, $T_{\rm K}$ = 600 K and $T_{\rm K}$ = 800 K, respectively. The black solid and dashed lines are thermal noise levels $\delta T^{\mathrm{N}}$ expected for an integration time of 100 h using SKA1-LOW and an integration time of 200 h using SKA2-LOW, respectively.
@@
The parameter (${m}_{\rm{WDM}}$, ${T}_{\rm{K}}$) distribution of the simulated data with an integration time of 100 h using SKA1-LOW (left panel) and 200 h using SKA2-LOW (right panel). In the left panel, black, blue and red correspond to the training set, verification set, and test set of the INF, respectively. The normalized parameter ($\tilde{m}_{\rm{WDM}}$, $\tilde{T}_{\rm{K}}$) axes are also presented. The test set of the GNF is the untrained part corresponding to the black data set. In the right panel, black and blue correspond to the training set and verification set of the INF, respectively. The test set of the GNF is the untrained part corresponding to the black data set.

### v2

@@ Methods
The relative error $\varepsilon$ of parameter constraints for the INF for different $T_{\rm K}$ and $m_{\rm WDM}$. a, c Present the temperature constrained precision of the $T_{\rm K}$ using SKA1-LOW and SKA2-LOW, respectively. b, d Present the constrained precision of $m_{\rm WDM}$ using SKA1-LOW and SKA2-LOW, respectively.
@@
The posterior distributions for $m_{\rm WDM}$ and $T_{\rm K}$ using different scales of 1D power spectra with the INF without considering thermal noise. The blue contour is based on the INF for $k$ in [$0.8~{}\rm Mpc^{-1}$, $3.5~{}\rm Mpc^{-1}$], the green contour is based on the INF for $k$ in [$5.5~{}\rm Mpc^{-1}$, $22.0~{}\rm Mpc^{-1}$], the orange contour is based on the INF for $k$ in [$35.2~{}\rm Mpc^{-1}$, $144.5~{}\rm Mpc^{-1}$] and the black contour is based on the INF for total scale. The panel presents the results when $m_{\rm WDM}=6$ keV and $T_{\rm K}=60$ K with an integration time of $100$ h using SKA1-LOW.
@@
Confidence interval distribution for distinguishing WDM from CDM under different $m_{\rm WDM}$ and $T_{\rm K}$. The color represents the confidence intervals, with blue dashed-dotted lines indicating the $1\sigma$ confidence level, dashed red lines indicating the $2\sigma$ confidence level, and solid black lines indicating the $3\sigma$ confidence level. a Presents the scenario with an SKA1-LOW integration time of $100$ h in the case of a weaker heating effect. b Presents the scenario with an SKA2-LOW integration time of $200$ h in the case of a stronger heating effect.
@@
Expected brightness temperature of the 21-cm forest. A total of 1 measurements of $10$ comoving megaparsec length segments along a line-of-sight neutral patch against 1 background light source with $S_{150}=10$ mJy yielded the expected value for the brightness temperature of 21-cm forest at $z=9$. a, c Present the brightness temperature for different $m_{\rm WDM}$ values in the WDM models. The blue, orange, and green curves correspond to $m_{\rm WDM}=3$ keV, $m_{\rm WDM}=6$ keV and $m_{\rm WDM}=9$ keV, respectively. $T_{\rm K}$ in (a) and (c) are 60 K and 600 K, respectively. b, d Present the brightness temperature corresponding to the different $T_{\rm K}$. In all cases, $m_{\rm WDM}$ is set to $6$ keV. The blue, orange, and green curves in (b) correspond to $T_{\rm K}=40$ K, $T_{\rm K}=60$ K and $T_{\rm K}=80$ K. The blue, orange, and green curves in (d) correspond to $T_{\rm K}=400$ K, $T_{\rm K}=600$ K and $T_{\rm K}=800$ K, respectively. The black solid and dashed lines are thermal noise levels $\delta T^{\mathrm{N}}$ expected for an integration time of $100$ h using SKA1-LOW and an integration time of $200$ h using SKA2-LOW, respectively.
@@
The parameter (${m}_{\rm{WDM}}$, ${T}_{\rm{K}}$) distribution of the simulated data. a Presents the distribution of the simulated data with an integration time of $100$ h using SKA1-LOW. b Presents the distribution of the simulated data with an integration time of $200$ h using SKA2-LOW. In both panels, black circle and blue block represent the training set and verification set of the GNF, respectively, while red rhombus corresponds to the testing set of the GNF (only in (a)). Green star represents randomly selected test data points. Both panels display the normalized parameter axes ($\tilde{m}_{\rm{WDM}}$, $\tilde{T}_{\rm{K}}$). For the INF, its testing set corresponds to the untrained portion of the black data set in both panels.
{sfigure}[Uncaptioned image]
[…]
{sfigure}[Uncaptioned image]
The mean values and $1\sigma$ distribution of the 1D power spectra generated by the GNF and linear interpolation methods, compared to simulated 1D power spectra. The data are derived from results of different $T_{\rm K}$ and $m_{\rm WDM}$, with an integration time of $100$ h using SKA1-LOW. a - c, d - f, g - i Present the 1D power spectra of 21-cm forest with $m_{\rm WDM}=4.5$ keV, $m_{\rm WDM}=6$ keV and $m_{\rm WDM}=7.5$ keV, respectively. The 1D power spectrum of 21-cm forest is presented in (a, d, g), (b, e, h) and (c, f, i) with $k=0.8~{}{\rm Mpc}^{-1}$ , $k=13.7~{}{\rm Mpc}^{-1}$ and $k=144.5~{}{\rm Mpc}^{-1}$, respectively. The blue shaded area represents the $1\sigma$ range of the 1D power spectra generated by the GNF, while the green shaded area represents the $1\sigma$ range of the 1D power spectra generated by linear interpolation. The orange line and bar indicate the mean value and $1\sigma$ error bar of the simulated power spectra.
{stable}
The key hyperparameters of the model algorithm. Hyperparameter GNF SKA1 GNF SKA2 INF SKA1 INF SKA2 Learning Rate $1\times 10^{-4}$ $1\times 10^{-4}$ $2\times 10^{-4}$ $2\times 10^{-4}$ Epochs Early Stopping Early Stopping $300$ $250$ Flow Steps $6$ $8$ $6$ $6$ Hidden Layer Size $256$ $256$ $128$ $256$ Transform Blocks $6$ $8$ $6$ $6$ Activation ReLU ReLU ReLU ReLU $L_{1}$ Regularization $3\times 10^{-6}$ $3\times 10^{-6}$ $0.0$ $0.0$ $L_{2}$ Regularization $1\times 10^{-4}$ $1\times 10^{-4}$ $0.0$ $0.0$ Dropout $0.0$ $0.0$ $0.0$ $0.0$ Batch Norm False False False False Bins $8$ $8$ $8$ $8$ Base Transform RQ-NSF (AR) RQ-NSF (AR) RQ-NSF (AR) RQ-NSF (AR)